# Learning Mean-Field Games through Mean-Field Actor-Critic Flow

### v1

@@ B.2 Convergence of the gap for value function
Next, we estimate each term in (B.43). Recall the definition
$V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=\sup_{(t^{\prime},x^{\prime})}\Big[V_{k}(t^{\prime},x^{\prime})-\frac{1}{2\iota}(|\widehat{t}_{k}-t^{\prime}|^{2}+|\widehat{x}_{k}-x^{\prime}|^{2})\Big].$
We denote $(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$ where this supremum is achieved. Similarly, we denote $(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime})$ where the infimum
$V_{k,\iota}^{*}(\widehat{s}_{k},\widehat{y}_{k})=\inf_{(s^{\prime},y^{\prime})}\Big[V_{k}^{*}(s^{\prime},y^{\prime})+\frac{1}{2\iota}(|\widehat{s}_{k}-s^{\prime}|^{2}+|\widehat{y}_{k}-y^{\prime}|^{2})\Big]$
is achieved. Then, by (A.48) in Lemma A.16
@@
where $C_{3}$ is recorded to determine $\iota$. We require that $\iota$ is small enough (cf. Figure 9) s.t. $C_{3}\,\iota\,R_{2}(2R_{2}^{2}+1)\leq 2r_{3}$. This implies
@@
Combining with (B.39) and (B.40) (and recall $2r_{3}\leq r_{2}$), we obtain
$\widehat{t}_{k}^{\prime},\widehat{s}_{k}^{\prime}\in[t_{-}+r_{2},t_{+}-r_{2}].$ (B.46)
By (A.52) in Lemma A.16, we have
$\partial_{t}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=\partial_{t}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),~~~~~\nabla_{x}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),~~~~\nabla_{x}^{2}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})\geq\nabla_{x}^{2}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$ (B.47)
@@
Then, we denote $(t_{k}^{\prime},x_{k}^{\prime})$ where the supremum for
$V_{k}^{\iota}(t_{k},x_{k})=\sup_{(t^{\prime},x^{\prime})}\Big[V_{k}(t^{\prime},x^{\prime})-\frac{1}{2\iota}(|t_{k}-t^{\prime}|^{2}+|x_{k}-x^{\prime}|^{2})\Big]$
is achieved. Similarly, we denote $(s_{k}^{\prime},y_{k}^{\prime})$ where the infimum for
$V_{k,\iota}^{*}(s_{k},y_{k})=\inf_{(s^{\prime},y^{\prime})}\Big[V_{k}^{*}(s^{\prime},y^{\prime})+\frac{1}{2\iota}(|s_{k}-s^{\prime}|^{2}+|y_{k}-y^{\prime}|^{2})\Big]$
is achieved. Lemma A.16 gives
$2(R_{2}-1)|t_{k}-t_{k}^{\prime}|+|x_{k}-x_{k}^{\prime}|\leq C_{3}\,\iota\,(R_{2}-1)(2(R_{2}-1)^{2}+1),\quad|x_{k}^{\prime}|\leq R_{2}-1$ (B.48)
$2(R_{2}-1)|s_{k}-s_{k}^{\prime}|+|y_{k}-y_{k}^{\prime}|\leq C_{3}\,\iota\,(R_{2}-1)(2(R_{2}-1)^{2}+1),\quad|y_{k}^{\prime}|\leq R_{2}-1$
$\partial_{t}V_{k}^{\iota}(t_{k},x_{k})=\partial_{t}V_{k}(t_{k}^{\prime},x_{k}^{\prime}),\quad~~\nabla_{x}V_{k}^{\iota}(t_{k},x_{k})=\nabla_{x}V_{k}(t_{k}^{\prime},x_{k}^{\prime})$ (B.49)
@@
Note that (B.48) implies
$|t_{k}-t_{k}^{\prime}|,|s_{k}-s_{k}^{\prime}|\leq r_{3}~~\Rightarrow~~t_{k}^{\prime},s_{k}^{\prime}\in[t_{-}+1.5r_{2},t_{+}-1.5r_{2}].$ (B.50)
Therefore, using (B.49) and the Hölder norm for $V_{k}$ (B.27), we have
@@
$\hskip 43.36243pt+C_{3}\,\iota\,R_{2}(2R_{2}^{2}+1)+r_{3}+C_{3}\,\iota\,(R_{2}-1)(2(R_{2}-1)^{2}+1)\Big]^{\zeta}$
$\leq C_{R_{2}}\big(\iota^{\frac{\zeta}{2}}+r_{3}^{\frac{\zeta}{2}}\big).$
where we used (B.40) (B.44), and (B.48) in the third inequality. Additionally, (B.49) and (B.27) also imply
@@
Next, we estimate the last term in (B.43). By definition (B.20), we have
$\partial_{t}\varphi(t,x,s,y)=-\gamma|x|^{l}_{l}+\frac{1}{\varepsilon}(t-s)-\frac{\lambda}{(t-t_{-})^{2}}.$
Therefore,
@@
where we used (B.39) and (B.41) in the third inequality, and (B.40) in the last inequality. Similarly, we can show
@@
For derivative in $x$, we have
$\nabla_{x}\varphi(t,x,s,y)=l\,\gamma(t_{+}-t+\Delta t)\,x^{l-1}+\frac{1}{\delta}(x-y),$
where we denote $x^{l-1}:=[x_{1}^{l-1},\ldots,x_{d}^{l-1}]^{\top}\in\mathbb{R}^{d}$ for simplicity. Therefore,
@@
$\leq l\,\gamma\left|\widehat{t}_{k}-t_{k}\right|\left|\widehat{x}_{k}^{l-1}\right|+l\,\gamma(t_{+}-\widehat{t}_{k}+\Delta t)\left|\widehat{x}_{k}^{l-1}-x_{k}^{l-1}\right|+\frac{1}{\delta}\left|(\widehat{x}_{k}-x_{k})-(\widehat{y}_{k}-y_{k})\right|$
$\leq l\,\gamma\,r_{3}\left|\widehat{x}_{k}\right|^{l-1}+2l\,\gamma\Delta t\,r_{3}+\frac{2r_{3}}{\delta}\leq\Big(l\,\gamma\,(R_{2}-1)^{l-1}+2l\,\gamma\Delta t+\frac{2}{\delta}\Big)r_{3},$
where we used (B.40) in the third inequality, and $|\widehat{x}_{k}|\leq R_{2}-1$ int he last inequality. Similarly, we can show
@@
Substituting (B.53), (B.54), and (B.59) into (B.43), we obtain
@@
where $C(R_{2},\gamma,\varepsilon,\delta,\lambda,\Delta t,r_{2})$ denotes a constant that depends on these parameters.
Step 5. In this step, we compute the critical point system for $\widehat{\Phi}_{k}$ and define a quantity $B_{k}$, which is the key to derive a contradiction. Since $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$ maximizes $\widehat{\Phi}_{k}$ and it lies the interior of $I_{t}\times B_{R_{2}}\times I_{t}\times B_{R_{2}}$, we have the first and second order necessary condition
@@
where a matrix inequality $A\leq B$ should be interpreted as BAB-A is positive semi-definite. Therefore
$\partial_{t}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=-\gamma|\widehat{x}_{k}|_{l}^{l}+\frac{1}{\varepsilon}(\widehat{t}_{k}-\widehat{s}_{k})-\frac{\lambda}{(\widehat{t}_{k}-t_{-})^{2}}+\mu(\widehat{t}_{k}-t_{k})-q$ (B.63)
@@
where $\nabla_{x}|x|_{l}^{l}:=(l|x_{1}|^{l-1}\textrm{sign}(x_{1}),\ldots,l|x_{d}|^{l-1}\textrm{sign}(x_{d}))^{\top}\in\mathbb{R}^{d}$. Similarly,
$\partial_{s}V_{k,\iota}^{*}(\widehat{t}_{k},\widehat{x}_{k})=-\gamma|\widehat{y}_{k}|_{l}^{l}+\frac{1}{\varepsilon}(\widehat{s}_{k}-\widehat{t}_{k})-\frac{\lambda}{(\widehat{s}_{k}-t_{-})^{2}}+\mu(\widehat{s}_{k}-s_{k})-\widehat{q}$ (B.65)
@@
Since $\nabla_{x}^{2}|x|_{l}^{l}=l(l-1)\,\textrm{diag}\big(|x_{1}|^{l-2},\ldots,|x_{d}|^{l-2}\big)$, $t_{+}-\widehat{t}_{k}+\Delta t\leq 2\Delta t$, and $\nabla_{x}\nabla_{y}|x|_{l}^{l}=\nabla_{x}\nabla_{y}|y|_{l}^{l}=0$, the last equation in (B.62) is simplified to
@@
$\widehat{D}^{k}_{x}=\textrm{diag}\big(|(\widehat{x}_{k})_{1}|^{l-2},\ldots,|(\widehat{x}_{k})_{d}|^{l-2}\big),~~\widehat{D}^{k}_{y}=\textrm{diag}\big(|(\widehat{y}_{k})_{1}|^{l-2},\ldots,|(\widehat{y}_{k})_{d}|^{l-2}\big).$ (B.68)
We define
$B_{k}:=\partial_{s}V_{k,\iota}^{*}(\widehat{s}_{k},\widehat{y}_{k})-\partial_{t}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=\partial_{s}V_{k}^{*}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime})-\partial_{t}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$ (B.69)
where the second equality is because of (B.47). By the optimality condition (B.63) and (B.65), we have
@@
If we use the PDE that $V_{k}$ and $V_{k}^{*}$ satisfies, we get
@@
We recall the local optimal $\alpha_{k}^{\diamond}$ is defined
$\alpha^{\diamond}_{k}(t,x):=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{n}}H(t,x,\mu_{t}^{k},\alpha,-\nabla_{x}V_{k}(t,x),-\nabla_{x}^{2}V_{k}(t,x)).$
We split (B.71) into two terms $B_{k}=(\text{I})+(\text{II})$, where
@@
$\quad-H\big(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},\alpha_{k}^{\diamond}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),-\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),-\nabla_{x}^{2}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})\big)$
and
@@
Next, we derive a local Lipschitz condition for the Hamiltonian in $\alpha$ and an upper bound for the local optimal $\alpha^{\diamond}$. For fixed $(t,x,\mu,p,P)\in[0,T]\times\mathbb{R}^{d}\times\mathcal{P}(\mathbb{R}^{d})\times\mathbb{R}^{d}\times\mathbb{R}^{d\times d}$ with $W_{2}(\mu,\delta)\leq K$, we temporally denote the map $\alpha\mapsto H(t,x,\mu,\alpha,p,P)$ as $H(\alpha)$. Then $H(\alpha)$ is $\lambda_{H}$-strongly concave. Let $\alpha^{\diamond}=\alpha^{\diamond}(t,x)$ be its maximizer (for given fixed $(\mu,p,P)$). For any $\alpha,\alpha^{\prime}\in\mathbb{R}^{d}$, we have
@@
$\leq K(1+|x|+C+|\alpha|\vee|\alpha^{\prime}|)|\alpha-\alpha^{\prime}|+K\,|\alpha-\alpha^{\prime}|\,|p|\leq C(1+|x|+|\alpha|\vee|\alpha^{\prime}|+|p|)\,|\alpha-\alpha^{\prime}|$
Additionally, the concavity of $H(\alpha)$ implies
$\left\langle\nabla_{\alpha}H(\alpha)-\nabla_{\alpha}H(\alpha^{\prime}),\,\alpha-\alpha^{\prime}\right\rangle\leq-\lambda_{H}|\alpha-\alpha^{\prime}|^{2}.$
Substitute $\alpha=0$ and $\alpha^{\prime}=\alpha^{\diamond}$ (note that $\nabla_{\alpha}H(\alpha^{\diamond})=0$), we get $\left\langle\nabla_{\alpha}H(0),\,\alpha^{\diamond}\right\rangle\geq\lambda_{H}|\alpha^{\diamond}|^{2}$, which implies $|\alpha^{\diamond}|\leq|\nabla_{\alpha}H(0)|/\lambda_{H}$. Therefore,
$\quad\left|\alpha^{\diamond}_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})\right|\leq\frac{1}{\lambda_{H}}\left(\left|\nabla_{\alpha}f(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},0)\right|+\left|\nabla_{\alpha}b(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},0)\right|\,\left|\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})\right|\right)$ (B.75)
@@
where we used Assumption 4.1 in the second inequality, and used (B.44) and (B.27) in the last inequality.
Next, we will show
@@
This time, we temporally denote the map $(t,x,\alpha)\mapsto H(t,x,\mu^{k}_{t},\alpha,-\nabla_{x}V_{k}(t,x),-\nabla_{x}^{2}V_{k}(t,x))$ by $H(t,x,\alpha)$. Given $(t,x),(t^{\prime},x^{\prime})$ (which will be replaced by $(\widehat{t}_{k},\widehat{x}_{k}),(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$ later), without loss of generality, we assume $H(t,x,\alpha^{\diamond}(t,x))\geq H(t^{\prime},x^{\prime},\alpha^{\diamond}(t^{\prime},x^{\prime}))$, then
@@
$\geq\frac{1}{2}\lambda_{H}\left|\alpha^{\diamond}(t,x)-\alpha^{\diamond}(t^{\prime},x^{\prime})\right|^{2}$
where the last inequality is because $H$ is $\lambda_{H}$-strongly concave in $\alpha$ and $\alpha^{\diamond}(t^{\prime},x^{\prime})$ maximize $H(t^{\prime},x^{\prime},\cdot)$. If $H(t^{\prime},x^{\prime},\alpha^{\diamond}(t^{\prime},x^{\prime}))>H(t,x,\alpha^{\diamond}(t,x))$, then we subtract $H(t,x,\alpha^{\diamond}(t^{\prime},x^{\prime}))$ (instead of $H(t^{\prime},x^{\prime},\alpha^{\diamond}(t,x))$) on both sides and get a similar inequality to (B.77). Next, we estimate the left-hand side of (B.77), where $(t,x)=(\widehat{t}_{k},\widehat{x}_{k})$, $(t^{\prime},x^{\prime})=(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$ satisfy (B.44), (B.45), and $V_{k}$ satisfies (B.27). By definition,
@@
$\quad+\left|\sigma(t,x,\mu^{k}_{t})-\sigma(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}})\right|\left|\nabla_{x}^{2}V_{k}(t,x)\right|+\left|\sigma(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}})\right|\left|\nabla_{x}^{2}V_{k}(t,x)-\nabla_{x}^{2}V_{k}(t^{\prime},x^{\prime})\right|$
Each term in (B.78) can be estimated as follows.

### v2

@@ B.2 Convergence of the gap for value function
We estimate each term on the right-hand side of (B.41). Denote by $(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$ where the supremum within $V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=\sup_{(t^{\prime},x^{\prime})}\Big[V_{k}(t^{\prime},x^{\prime})-\frac{1}{2\iota}(|\widehat{t}_{k}-t^{\prime}|^{2}+|\widehat{x}_{k}-x^{\prime}|^{2})\Big]$ is attained. Similarly, denote by $(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime})$ where the infimum within $V_{k,\iota}^{*}(\widehat{s}_{k},\widehat{y}_{k})=\inf_{(s^{\prime},y^{\prime})}\Big[V_{k}^{*}(s^{\prime},y^{\prime})+\frac{1}{2\iota}(|\widehat{s}_{k}-s^{\prime}|^{2}+|\widehat{y}_{k}-y^{\prime}|^{2})\Big]$ is attained. The existence of $(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$ and $(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime})$ follows from Lemma A.16. By (A.48),
@@
where $C_{3}$ is recorded for later specification of $\iota$. Set $\iota$ to be small enough (cf. Figure 9) such that $C_{3}\,\iota\,R_{2}(2R_{2}^{2}+1)\leq 2r_{3}$. This implies
@@
Combining with (B.37) and (B.38) (and recall $2r_{3}\leq r_{2}$) yields $\widehat{t}_{k}^{\prime},\widehat{s}_{k}^{\prime}\in[t_{-}+r_{2},t_{+}-r_{2}]$. By (A.52) in Lemma A.16,
$\partial_{t}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=\partial_{t}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),~~~~~\nabla_{x}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),~~~~\nabla_{x}^{2}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})\geq\nabla_{x}^{2}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),$ (B.44)
@@
Denote by $(t_{k}^{\prime},x_{k}^{\prime})$ where the supremum within $V_{k}^{\iota}(t_{k},x_{k})=\sup_{(t^{\prime},x^{\prime})}\Big[V_{k}(t^{\prime},x^{\prime})-\frac{1}{2\iota}(|t_{k}-t^{\prime}|^{2}+|x_{k}-x^{\prime}|^{2})\Big]$ is attained. Similarly, denote by $(s_{k}^{\prime},y_{k}^{\prime})$ where the infimum within $V_{k,\iota}^{*}(s_{k},y_{k})=\inf_{(s^{\prime},y^{\prime})}\Big[V_{k}^{*}(s^{\prime},y^{\prime})+\frac{1}{2\iota}(|s_{k}-s^{\prime}|^{2}+|y_{k}-y^{\prime}|^{2})\Big]$ is attained. By Lemma A.16,
$2(R_{2}-1)|t_{k}-t_{k}^{\prime}|+|x_{k}-x_{k}^{\prime}|\leq C_{3}\,\iota\,(R_{2}-1)(2(R_{2}-1)^{2}+1),\quad|x_{k}^{\prime}|\leq R_{2}-1,$ (B.45)
$2(R_{2}-1)|s_{k}-s_{k}^{\prime}|+|y_{k}-y_{k}^{\prime}|\leq C_{3}\,\iota\,(R_{2}-1)(2(R_{2}-1)^{2}+1),\quad|y_{k}^{\prime}|\leq R_{2}-1,$
$\partial_{t}V_{k}^{\iota}(t_{k},x_{k})=\partial_{t}V_{k}(t_{k}^{\prime},x_{k}^{\prime}),\quad~~\nabla_{x}V_{k}^{\iota}(t_{k},x_{k})=\nabla_{x}V_{k}(t_{k}^{\prime},x_{k}^{\prime}),$ (B.46)
@@
Note that (B.45) implies $|t_{k}-t_{k}^{\prime}|,|s_{k}-s_{k}^{\prime}|\leq r_{3}$, which leads to $t_{k}^{\prime},s_{k}^{\prime}\in[t_{-}+1.5r_{2},t_{+}-1.5r_{2}]$. Therefore, using (B.44), (B.46) and the Hölder norm bound for $V_{k}$ (B.26),
@@
$\hskip 43.36243pt+C_{3}\,\iota\,R_{2}(2R_{2}^{2}+1)+r_{3}+C_{3}\,\iota\,(R_{2}-1)(2(R_{2}-1)^{2}+1)\Big]^{\zeta}\leq C_{R_{2}}\big(\iota^{\frac{\zeta}{2}}+r_{3}^{\frac{\zeta}{2}}\big).$
where the third inequality is from (B.38) (B.42), and (B.45). Additionally, (B.46) and (B.26) imply
@@
We estimate the last term in (B.41). By (B.20), $\partial_{t}\varphi(t,x,s,y)=-\gamma|x|^{l}_{l}+\frac{1}{\varepsilon}(t-s)-\frac{\lambda}{(t-t_{-})^{2}}.$ Therefore,
@@
where we use (B.37), (B.38) and (B.39). Similarly,
@@
For derivatives in $x$, $\nabla_{x}\varphi(t,x,s,y)=l\,\gamma(t_{+}-t+\Delta t)\,x^{l-1}+\frac{1}{\delta}(x-y),$ where $x^{l-1}$ denotes the component-wise power of $x$. Therefore,
@@
$\leq l\,\gamma\left|\widehat{t}_{k}-t_{k}\right|\left|x_{k}^{l-1}\right|+l\,\gamma(t_{+}-\widehat{t}_{k}+\Delta t)\left|\widehat{x}_{k}^{l-1}-x_{k}^{l-1}\right|+\frac{1}{\delta}\left|(\widehat{x}_{k}-x_{k})-(\widehat{y}_{k}-y_{k})\right|$
$\leq l\,\gamma\,r_{3}\left|x_{k}\right|^{l-1}+2l\,\gamma\Delta t\,r_{3}+\frac{2r_{3}}{\delta}\leq\Big(l\,\gamma\,(R_{2}-1)^{l-1}+2l\,\gamma\Delta t+\frac{2}{\delta}\Big)r_{3}.$
Similarly,
@@
Substituting (B.49), (B.50), and (B.55) into (B.41) yields
@@
where $C(R_{2},\gamma,\varepsilon,\delta,\lambda,\Delta t,r_{2})$ denotes a constant that depends on those parameters.
Step 5. In this step, we compute the critical point system for $\widehat{\Phi}_{k}$ and define a quantity $B_{k}$, which is the key to deriving a contradiction. Since $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$ maximizes $\widehat{\Phi}_{k}$ in the interior of $I_{t}\times B_{R_{2}}\times I_{t}\times B_{R_{2}}$, the first and second order necessary conditions provide
@@
where matrix inequalities are in positive semi-definite sense. Therefore
$\partial_{t}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=-\gamma|\widehat{x}_{k}|_{l}^{l}+\frac{1}{\varepsilon}(\widehat{t}_{k}-\widehat{s}_{k})-\frac{\lambda}{(\widehat{t}_{k}-t_{-})^{2}}+\mu(\widehat{t}_{k}-t_{k})-q,$ (B.59)
@@
where $\nabla_{x}|x|_{l}^{l}:=lx^{l-1}\in\mathbb{R}^{d}$, where the power applies component-wise. Similarly,
$\partial_{s}V_{k,\iota}^{*}(\widehat{t}_{k},\widehat{x}_{k})=-\gamma|\widehat{y}_{k}|_{l}^{l}+\frac{1}{\varepsilon}(\widehat{s}_{k}-\widehat{t}_{k})-\frac{\lambda}{(\widehat{s}_{k}-t_{-})^{2}}+\mu(\widehat{s}_{k}-s_{k})-\widehat{q}$, (B.61)
@@
Since $\nabla_{x}^{2}|x|_{l}^{l}=l(l-1)\,\textrm{diag}\big(x^{l-2}\big)$, $t_{+}-\widehat{t}_{k}+\Delta t\leq 2\Delta t$, the last equation in (B.58) simplifies to
@@
$\widehat{D}^{k}_{x}=\textrm{diag}\big(\widehat{x}_{k}^{l-2}\big),~~\widehat{D}^{k}_{y}=\textrm{diag}\big(\widehat{y}_{k}^{l-2}\big).$ (B.64)
Define $B_{k}:=\partial_{s}V_{k,\iota}^{*}(\widehat{s}_{k},\widehat{y}_{k})-\partial_{t}V_{k}^{\iota}(\widehat{t}_{k},\widehat{x}_{k})=\partial_{s}V_{k}^{*}(\widehat{s}_{k}^{\prime},\widehat{y}_{k}^{\prime})-\partial_{t}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$, where the second equality follows from (B.44). The optimality conditions (B.59) and (B.61) imply
@@
Using the PDEs that characterize $V_{k}$ and $V_{k}^{*}$, we get
@@
Recall that $\alpha^{\diamond}_{k}(t,x):=\operatorname*{arg\,max}_{a\in\mathbb{R}^{n}}H(t,x,\mu_{t}^{k},a,-\nabla_{x}V_{k}(t,x),-\nabla_{x}^{2}V_{k}(t,x)).$ We split (B.66) into two terms $B_{k}=(\text{I})+(\text{II})$, where
@@
$\quad-H\big(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},\alpha_{k}^{\diamond}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),-\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime}),-\nabla_{x}^{2}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})\big),$
@@
Next, we prove a local Lipschitz condition of the Hamiltonian in $\alpha$ and establish a bound for the local optimal control $\alpha^{\diamond}$. For fixed $(t,x,\mu,p,P)\in[0,T]\times\mathbb{R}^{d}\times\mathcal{P}(\mathbb{R}^{d})\times\mathbb{R}^{d}\times\mathbb{R}^{d\times d}$ with $W_{2}(\mu,\delta_{0})\leq K$, we temporarily denote the $\lambda_{H}$-strongly concave mapping $\alpha\mapsto H(t,x,\mu,\alpha,p,P)$ by $H(\alpha)$. Let $\alpha^{\diamond}$ denote its maximizer for given fixed $(\mu,p,P)$. For any $\alpha,\alpha^{\prime}\in\mathbb{R}^{n}$,
@@
$\leq K(1+|x|+K+|\alpha|\vee|\alpha^{\prime}|)|\alpha-\alpha^{\prime}|+K\,|\alpha-\alpha^{\prime}|\,|p|\leq C(1+|x|+|\alpha|\vee|\alpha^{\prime}|+|p|)\,|\alpha-\alpha^{\prime}|.$
By the concavity of $H(\alpha)$, $\left\langle\nabla_{\alpha}H(\alpha)-\nabla_{\alpha}H(\alpha^{\prime}),\,\alpha-\alpha^{\prime}\right\rangle\leq-\lambda_{H}|\alpha-\alpha^{\prime}|^{2}.$ Substituting $\alpha=0$, $\alpha^{\prime}=\alpha^{\diamond}$ and using $\nabla_{\alpha}H(\alpha^{\diamond})=0$, we get $\left\langle\nabla_{\alpha}H(0),\,\alpha^{\diamond}\right\rangle\geq\lambda_{H}|\alpha^{\diamond}|^{2}$, which implies $|\alpha^{\diamond}|\leq|\nabla_{\alpha}H(0)|/\lambda_{H}$. Therefore,
$\left|\alpha^{\diamond}_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})\right|$ $\leq\frac{1}{\lambda_{H}}\left(\left|\nabla_{\alpha}f(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},0)\right|+\left|\nabla_{\alpha}b(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime},\mu^{k}_{\widehat{t}_{k}^{\prime}},0)\right|\,\left|\nabla_{x}V_{k}(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})\right|\right)$ (B.70)
@@
where we use Assumption 4.1, (B.26) and (B.42).
As a next step, we show that
@@
This time, we temporarily denote the mapping $(t,x,\alpha)\mapsto H(t,x,\mu^{k}_{t},\alpha,-\nabla_{x}V_{k}(t,x),-\nabla_{x}^{2}V_{k}(t,x))$ by $H(t,x,\alpha)$. For any pair of tuples $(t,x),(t^{\prime},x^{\prime})$ (later evaluated at $(\widehat{t}_{k},\widehat{x}_{k}),(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$), without loss of generality, we assume $H(t,x,\alpha^{\diamond}(t,x))\geq H(t^{\prime},x^{\prime},\alpha^{\diamond}(t^{\prime},x^{\prime}))$, which implies
@@
$\geq\frac{1}{2}\lambda_{H}\left|\alpha^{\diamond}(t,x)-\alpha^{\diamond}(t^{\prime},x^{\prime})\right|^{2},$
where the last inequality follows from the fact that $H$ is $\lambda_{H}$-strongly concave in $\alpha$ and that $\alpha^{\diamond}(t^{\prime},x^{\prime})$ maximizes $H(t^{\prime},x^{\prime},\cdot)$. We remark that, if the converse $H(t^{\prime},x^{\prime},\alpha^{\diamond}(t^{\prime},x^{\prime}))>H(t,x,\alpha^{\diamond}(t,x))$ holds, subtracting $H(t,x,\alpha^{\diamond}(t^{\prime},x^{\prime}))$ (instead of $H(t^{\prime},x^{\prime},\alpha^{\diamond}(t,x))$) on both sides yields a similar inequality to (B.72). We estimate the left-hand side of (B.72), evaluated at $(t,x)=(\widehat{t}_{k},\widehat{x}_{k})$, $(t^{\prime},x^{\prime})=(\widehat{t}_{k}^{\prime},\widehat{x}_{k}^{\prime})$, with $V_{k}$ satisfying (B.26).
@@
$\quad+\left|\sigma(t,x,\mu^{k}_{t})-\sigma(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}})\right|\left|\nabla_{x}^{2}V_{k}(t,x)\right|+\left|\sigma(t^{\prime},x^{\prime},\mu^{k}_{t^{\prime}})\right|\left|\nabla_{x}^{2}V_{k}(t,x)-\nabla_{x}^{2}V_{k}(t^{\prime},x^{\prime})\right|.$
Based on Assumption 4.1 and Assumption 4.3, each term in (B.73) can be estimated as follows: